# arXiv:2603.01338v1[math.AP] 02 Mar 2026

### v1

@@ Abstract.
We study the scattering problem for the three dimensional Zakharov-Kuznetsov equation in the framework of the final state problem. We construct a global solution to the Zakharov-Kuznetsov equation which scatters to a given free solution.
@@ 1. Introduction
This paper is concerned with the scattering problem for the Zakharov-Kuznetsov equation in three dimensions:
@@
where $u:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$ is an unknown function, $x=(x_{1},x_{2},x_{3})$ and $\Delta$ is Laplacian on $\mathbb{R}^{3}$. Equation (1.1) was derived by Zakharov-Kuznetsov [27] to describe unidirectional wave propagation in a magnetized plasma. See also Laedke-Spatschek [18] for the derivation of (1.1) from the basic hydrodynamical equations, and Lannes-Linares-Saut [19] for the rigorous justification of (1.1) from the Euler-Poisson system for a uniformly magnetized media.
Equation (1.1) has the conservation of mass
(1.2) $M[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}u^{2}dx=M[u](0)$
for any $t\in\mathbb{R}$, and the conservation of energy
(1.3) $E[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}|\nabla u|^{2}dx-\frac{1}{6}\int_{\mathbb{R}^{3}}u^{3}dxdy=E[u](0)$
for any $t\in\mathbb{R}$.
@@
has been studied from the point of view of well-posedness [5, 10, 12, 13, 15, 20, 21, 24], and stability of soliton [3, 4, 6, 17, 26] etc. Concerning the scattering problem for (1.4), from the fact that the solution of the linear equation associated with (1.4) decays like $O(t^{-d/2})$ in $L^{\infty}$ as $t\to\infty$ (see [14, Theorem 3.2] for instance), and from the point of view of the linear scattering theory (see [23] for instance), we expect that if $d\geq 3$, then (at least small) solution to (1.4) scatters to the free solution. Herr-Kinoshita [12] proved the small data scattering for the initial value problem of (1.4) with $d\geq 5$ in the scaling critical Sobolev space. Furthermore, they proved the scattering for (1.4) with $d=4$ when the initial data is small and radial in the last (d1)(d-1) variables. For two dimensional case $d=2$, the author [25] proved the existence of small global solutions to (1.4) which scatters to a given free solution. See also Farah-Linares-Pastor [7], Anjolras [1] Correia-Kinoshita [2] for the scattering results on (1.4) with $d=2$ and power type nonlinearity with degree higher than two.
In this paper we consider the scattering problem for (1.4) with $d=3$ in the framework of the final state problem. To state the our main theorem, we introduce several notation. For $0<\delta<1$, we define a semi-normed space $(X_{\delta},\|\cdot\|_{X_{\delta}}$) by
@@
In the definition of the function space $X_{\delta}$, we required that $f\in{{\mathcal{S}}}$ for the simplicity of the argument. Although this requirement can be relaxed (for example, by the density argument), we do not discuss about it in this paper.
We now give the outline of the proof of Theorem 1.1. For given final data $u_{+}\in X_{\delta}$ with $0<\delta<1$, we introduce a new unknown function
@@
with suitable polynomials $p,q,r$ and $s$, where $\sum p(\xi)q(\eta)$ denotes finite sum of the functions of form $p(\xi)q(\eta)$, see Lemma 3.2 below for detail. Combining (1.11) with integration by parts both in space and time variables, we derive the time decay of (1) in $L^{2}$. Note that the null structure of the nonlinearity is employed in various contexts to study global dynamics of nonlinear PDEs since the pioneering work by Klainerman [16]. We note that this approach was also used in [25] to prove the existence of small global solution to (1.1) with $d=2$ which scatters to the given free solution. Furthermore, in [25], the author transformed (1.4) into the equation which is symmetric with respect to $x_{1}$ and $x_{2}$. Thanks to this transform, the problem became more transparent in two dimensional case. On the other hand, the similar transform is not known for (1.4) with $d\neq 2$. Therefore derivation of the key identity (1.11) for $d=3$ is more complicated compared to two dimensional case. Once we obtain $L^{2}$ estimate for $u_{2}$, we have an existence of solution to (1.9) by the compactness argument.
@@
The outline of the paper is as follows. In Section 2, we give the decay and Strichartz estimates for the linearized equation of (1.9). In Section 3, we derive the key bilinear dispersive estimates. Finally, in Section 4, we prove Theorem 1.1.
@@ 3. Bilinear Dispersive Estimates
Simple calculation yields
@@
(3.6) $\psi_{\text{time}}(\xi,\eta)$ $=$ $\left(A_{0,0}(\eta)+\xi_{1}A_{0,1}(\eta)+\xi_{2}A_{0,2}(\eta)+\xi_{3}A_{0,3}(\eta)\right)\phi$
$\ \ \psi_{\text{space}}(\xi,\eta)$ $=$ $\left(A_{1,0}(\eta)+\xi_{1}A_{1,1}(\eta)+\xi_{2}A_{1,2}(\eta)+\xi_{3}A_{1,3}(\eta)\right)\partial_{\eta_{1}}\phi$
$+\left(A_{2,0}(\eta)+\xi_{1}A_{2,1}(\eta)+\xi_{2}A_{2,2}(\eta)\right$.
@@
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}A_{2,3}(\eta)+\xi_{3}^{2}A_{2,4}(\eta)\right)\partial_{\eta_{2}}\phi$
$+\left(A_{3,0}(\eta)+\xi_{1}A_{3,1}(\eta)+\xi_{3}A_{3,2}(\eta)\right$.
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}A_{3,3}(\eta)+\xi_{2}^{2}A_{3,4}(\eta)\right)\partial_{\eta_{3}}\phi,$
where
$A_{0,0}(\eta)$ $=$ $\frac{1}{3p(\eta)}\{9\eta_{1}^{4}+30\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+5(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{0,1}(\eta)$ $=$ $-\frac{4}{p(\eta)}\eta_{1}(\eta_{2}^{2}+\eta_{3}^{2}),\quad A_{0,2}(\eta)=-\frac{2}{3p(\eta)}\eta_{2}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{0,3}(\eta)$ $=$ $-\frac{2}{3p(\eta)}\eta_{3}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{1,0}(\eta)$ $=$ $-\frac{1}{3p(\eta)}\eta_{1}\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{1,1}(\eta)$ $=$ $\frac{4}{p(\eta)}\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2}),\quad A_{1,2}(\eta)=\frac{2}{3p(\eta)}\eta_{1}\eta_{2}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{1,3}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{1}\eta_{3}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{i,0}(\eta)$ $=$ $-\frac{4}{3p(\eta)}\eta_{i}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad A_{i,1}(\eta)=\frac{4}{p(\eta)}\eta_{1}\eta_{i}(\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{i,2}(\eta)$ $=$ $\frac{2}{3p(\eta)}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad A_{i,3}(\eta)=-\frac{2}{3p(\eta)}\eta_{5-i}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),$
$A_{i,4}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{i}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\qquad i=2,3$
with $p(\eta)=(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}$.
Proof of Lemma 3.2..
By (3.3), we see
(3.8) $\partial_{\eta_{1}}\phi(\xi,\eta)$ $=$ $3\xi_{1}^{2}+\xi_{2}^{2}+\xi_{3}^{2}-6\xi_{1}\eta_{1}-2\xi_{2}\eta_{2}-2\xi_{3}\eta_{3},$
(3.9) $\partial_{\eta_{2}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{2}-2\xi_{1}\eta_{2}-2\eta_{1}\xi_{2},$
(3.10) $\partial_{\eta_{3}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{3}-2\xi_{1}\eta_{3}-2\eta_{1}\xi_{3}.$
Hence, we have
Therefore,
(3.11) $\eta_{2}\xi_{2}+\eta_{3}\xi_{3}=-\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{2\eta_{1}}\xi_{1}+N_{1},$
where $N_{1}=N_{1}(\xi,\eta)$ is given by
(3.12) $N_{1}=\frac{1}{2\eta_{1}}\left(\phi(\xi,\eta)-\eta\cdot\nabla_{\eta}\phi(\xi,\eta)\right).$
[…]
Therefore,
@@
(3.13) $\eta_{3}\xi_{2}-\eta_{2}\xi_{3}$ $=$ $\frac{1}{2\eta_{1}}\left\{(\xi_{3}-\eta_{3})\partial_{\eta_{2}}\phi-(\xi_{2}-\eta_{2})\partial_{\eta_{3}}\phi\right\}.$
By (3.11) and (3.13),
where $N_{2}=N_{2}(\xi,\eta)$ is given by
Hence
(3.15) $\xi_{2}^{2}+\xi_{3}^{2}=\frac{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}+N_{2}(\xi,\eta).$
Substituting (3.11) and (3.15) into (3.8), we have
$\partial_{\eta_{1}}\phi(\xi,\eta)$
Therefore
$=$ $\partial_{\eta_{1}}\phi+2N_{1}-N_{2}.$
[…]
Hence by (3.11), we obtain
Therefore,
By ${(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})/\eta_{1}^{2}\times}$ $\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\}/\{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})\}\times(3)$, we have
where $N_{3}=N_{3}(\xi,\eta)$ is given by
Hence we obtain
(3.19) $\xi_{1}=\frac{4\eta_{1}^{3}(\eta_{2}^{2}+\eta_{3}^{2})}{3(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}}N_{3}.$
Substituting (3.12), (LABEL:n2), (LABEL:n3) into (3.19), we have (3.5). ∎
@@
By using (3.5), we split $I(f,g)$ defined by (3.4) into several terms:
@@
We first evaluate $I_{\text{time}}(f,g)$. We treat the second term of $\psi_{\text{time}}$, i.e., $\xi_{1}A_{0,1}\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{time},2}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}A_{0,1}(\eta)\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
@@
$I_{\text{time},2}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\partial_{\tau}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}A_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $i\limsup_{T\to\infty}\xi_{1}\int_{\mathbb{R}^{3}}e^{-iT\phi(\xi,\eta)}A_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta$
$-i\xi_{1}\int_{\mathbb{R}^{3}}e^{-it\phi(\xi,\eta)}A_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta.$
@@
$\lesssim$ $\limsup_{T\to\infty}\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(T)f](\xi-\eta){{\mathcal{F}}}\left[\frac{\partial_{x_{1}}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(T)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[\frac{\partial_{x_{1}}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(t)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$\lesssim$ $\limsup_{T\to\infty}\|\partial_{x_{1}}V(T)f\|_{L_{x}^{\infty}}\left\|\frac{\partial_{x_{1}}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(T)g\right\|_{L_{x}^{2}}$
$+\limsup_{T\to\infty}\|V(T)f\|_{L_{x}^{\infty}}\left\|\frac{\partial_{x_{1}}^{2}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(T)g\right\|_{L_{x}^{2}}$
$+\|\partial_{x_{1}}V(t)f\|_{L_{x}^{\infty}}\left\|\frac{\partial_{x_{1}}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(t)g\right\|_{L_{x}^{2}}$
$+\|V(t)f\|_{L_{x}^{\infty}}\left\|\frac{\partial_{x_{1}}^{2}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(t)g\right\|_{L_{x}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the boundedness of $\Delta^{-1}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})$ in $L^{2}$, we have
@@
Next we evaluate $I_{\text{space}}(f,g)$. We treat the second term of $\psi_{\text{space}}$, i.e., $\xi_{1}A_{1,1}(\eta)\partial_{\eta_{1}}\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{space,2}}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}A_{1,1}(\eta)\partial_{\eta_{1}}\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
By an integration by parts in $\eta$, we have
$I_{\text{space,2}}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}\partial_{\eta_{1}}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}A_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}\partial_{\eta_{1}}A_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$+i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}A_{1,1}(\eta)\partial_{\eta_{1}}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}A_{1,1}(\eta)\widehat{f}(\xi-\eta)\partial_{\eta_{1}}\widehat{g}(\eta)d\eta d\tau.$

### v2

# Scattering problem for Zakharov-Kuznetsov equation in three space dimensions

@@ Abstract.
This paper is a continuation of our previous study [25] on the scattering problem for the Zakharov-Kuznetsov equation (ZK). When the space dimension is three, we construct a global solution to (ZK) which scatters to a given free solution without smallness assumption on the asymptotic states.
@@ 1. Introduction
This paper is a continuation of our previous study [25] on the scattering problem for the Zakharov-Kuznetsov equation. In this paper we focus on the Zakharov-Kuznetsov equation in three dimensions:
@@
where $u:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$ is an unknown function, $x=(x_{1},x_{2},x_{3})$ and $\Delta$ is Laplacian on $\mathbb{R}^{3}$. Equation (1.1) was derived by Zakharov-Kuznetsov [27] to describe unidirectional wave propagation in a magnetized plasma. Note that Laedke-Spatschek [18] derived (1.1) from the basic hydrodynamical equations. Furthermore, Lannes-Linares-Saut [19] gave the rigorous justification of (1.1) from the Euler-Poisson system for a uniformly magnetized media.
Equation (1.1) has the conservation of mass : for any $t\in\mathbb{R}$,
(1.2) $M[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}u(t,x)^{2}dx=M[u](0),$
and the conservation of energy : for any $t\in\mathbb{R}$,
(1.3) $\ \ \ E[u](t):=\frac{1}{2}\int_{\mathbb{R}^{3}}|(\nabla u)(t,x)|^{2}dx-\frac{1}{6}\int_{\mathbb{R}^{3}}u(t,x)^{3}dxdy=E[u](0).$
@@
has been studied from the point of view of well-posedness [5, 10, 12, 13, 15, 20, 21, 24], and stability of soliton [3, 4, 6, 17, 26] etc. Concerning the scattering problem for (1.4), from the fact that the solution of the linear equation associated with (1.4) decays like $O(t^{-d/2})$ in $L^{\infty}$ as $t\to\infty$ (see [14, Theorem 3.2] for instance), and from the point of view of the linear scattering theory (see [23] for instance), we expect that if $d\geq 3$, then (at least small) solution to (1.4) scatters to the free solution. Herr-Kinoshita [12] proved the small data scattering for the initial value problem of (1.4) with $d\geq 5$ in the scaling critical Sobolev space. Furthermore, they proved the scattering for (1.4) with $d=4$ when the initial data is small and radial in the last (d1)(d-1) variables. For two dimensional case $d=2$, the author [25] proved the existence of small global solutions to (1.4) which scatters to a given free solution. See also Farah-Linares-Pastor [7], Anjolras [1], and Correia-Kinoshita [2] for the scattering results on (1.4) with $d=2$ and power type nonlinearity with degree higher than two.
In this paper we consider the scattering problem for (1.4) with physically important case $d=3$ in the framework of the final state problem. To state the our main theorem, we introduce several notation. For $0<\delta<1$, we define a semi-normed space $(X_{\delta},\|\cdot\|_{X_{\delta}}$) by
@@
In the definition of the function space $X_{\delta}$, we required that $f\in{{\mathcal{S}}}$ for the simplicity of the argument. This requirement can be relaxed by the density argument for example. However, we do not discuss about it in this paper.
We now give the outline of the proof of Theorem 1.1. As in [25], for given final data $u_{+}\in X_{\delta}$, we introduce a new unknown function
@@
with suitable polynomials $p,r$ and rational functions $q,s$, see Lemma 3.2 below for detail. Combining (1.11) with integration by parts both in $\tau$ and $\eta$, we derive the time decay of (1) in $L^{2}$. Note that the null structure of the nonlinearity is employed in various contexts to study global dynamics of nonlinear PDEs since the pioneering work by Klainerman [16]. We note that this approach was also used in [25] to prove the existence of small global solution to (1.1) with $d=2$ which scatters to the given free solution. Furthermore, in [25], the author transformed (1.4) into the equation which is symmetric with respect to $x_{1}$ and $x_{2}$. Thanks to this transform, the problem became more transparent in two dimensional case. On the other hand, the similar transform is not known for (1.4) with $d\neq 2$. Therefore derivation of the key identity (1.11) for $d=3$ is more complicated compared to two dimensional case. Once we obtain $L^{2}$ estimate for $u_{2}$, we have an existence of solution to (1.9) by the compactness argument.
@@
The outline of the paper is as follows. In Section 2, we give the decay and Strichartz estimates for the linearized equation of (1.9). In Section 3, we derive the key bilinear dispersive estimates. Finally, in Section 4, we prove Theorem 1.1. In Appendix, we give the proof of Lemma 3.2.
@@ 3. Bilinear Dispersive Estimates
To prove Proposition 3.1, we need to do some preparation. Simple calculation yields
@@
(3.6) $\psi_{\text{time}}(\xi,\eta)$ $=$ $\left(A_{0}(\eta)+\xi_{1}B_{0,1}(\eta)+\xi_{2}B_{0,2}(\eta)+\xi_{3}B_{0,3}(\eta)\right)\phi$
$\ \ \ \psi_{\text{space}}(\xi,\eta)$ $=$ $\left(A_{1}(\eta)+\xi_{1}B_{1,1}(\eta)+\xi_{2}B_{1,2}(\eta)+\xi_{3}B_{1,3}(\eta)\right)\partial_{\eta_{1}}\phi$
$+\left(A_{2}(\eta)+\xi_{1}B_{2,1}(\eta)+\xi_{2}B_{2,2}(\eta)\right$.
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}C_{2,1}(\eta)+\xi_{3}^{2}C_{2,2}(\eta)\right)\partial_{\eta_{2}}\phi$
$+\left(A_{3}(\eta)+\xi_{1}B_{3,1}(\eta)+\xi_{3}B_{3,2}(\eta)\right$.
$\qquad\qquad\qquad\qquad\left.+\xi_{2}\xi_{3}C_{3,1}(\eta)+\xi_{2}^{2}C_{3,2}(\eta)\right)\partial_{\eta_{3}}\phi,$
where $A_{j}$, $B_{j,k}$ and $C_{j,k}$ are rational functions in $\eta$ satisfying
$|A_{0}(\eta)|$ $\lesssim$ $(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2},$
@@
$|\partial_{\eta_{j}}^{i}A_{j}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{3}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\qquad(i=0),\\ &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\qquad(i=1)\end{aligned}\right$.
for $j=1,2,3$,
$|B_{0,k}(\eta)|$ $\lesssim$ $(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}$
for $k=1,2,3$,
@@
$|\partial_{\eta_{j}}^{i}B_{j,k}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\quad(i=0),\\ &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{1}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{3}{2}}|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\quad(i=1)\end{aligned}\right$.
for (j,k)=(1,1),(1,2),(1,3),(2,1),(2,2),(3,1),(3,2)(j,k)=(1,1),(1,2),(1,3),(2,1),(2,2),(3,1),(3,2),
$|\partial_{\eta_{j}}^{i}C_{j,k}(\eta)|$ $\lesssim$ $\left\{\begin{aligned} &(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{\frac{1}{2}}(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\quad(i=0),\\ &(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{-2}\\ &+(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})|3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2}|^{-3}\quad(i=1),\end{aligned}\right$.
for $(j,k)=(2,1),(2,2),(3,1),(3,2)$.
We shall prove Lemma 3.2 in Appendix.
@@
By using (3.5), we split $I(f,g)$ defined by (3.4) into the following two terms:
@@
We first evaluate $I_{\text{time}}(f,g)$. We treat the second term of $\psi_{\text{time}}$, i.e., $\xi_{1}B_{0,1}(\eta)\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{time},2}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}B_{0,1}(\eta)\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
@@
$I_{\text{time},2}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\partial_{\tau}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $i\limsup_{T\to\infty}\xi_{1}\int_{\mathbb{R}^{3}}e^{-iT\phi(\xi,\eta)}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta$
$-i\xi_{1}\int_{\mathbb{R}^{3}}e^{-it\phi(\xi,\eta)}B_{0,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta.$
@@
$\lesssim$ $\limsup_{T\to\infty}\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(T)f](\xi-\eta){{\mathcal{F}}}\left[B_{0,1}(-i\nabla)V(T)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{\mathbb{R}^{3}}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[B_{0,1}(-i\nabla)V(t)g\right](\eta)d\eta\right\|_{L_{\xi}^{2}}$
$\lesssim$ $\limsup_{T\to\infty}\|\partial_{x_{1}}V(T)f\|_{L_{x}^{\infty}}\left\|B_{0,1}(-i\nabla)V(T)g\right\|_{L_{x}^{2}}$
$+\limsup_{T\to\infty}\|V(T)f\|_{L_{x}^{\infty}}\left\|\partial_{x_{1}}B_{0,1}(-i\nabla)V(T)g\right\|_{L_{x}^{2}}$
$+\|\partial_{x_{1}}V(t)f\|_{L_{x}^{\infty}}\left\|B_{0,1}(-i\nabla)V(t)g\right\|_{L_{x}^{2}}$
$+\|V(t)f\|_{L_{x}^{\infty}}\left\|\partial_{x_{1}}B_{0,1}(-i\nabla)V(t)g\right\|_{L_{x}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the inequality for $B_{0,1}$ in Lemma 3.2, we have
@@
Next we evaluate $I_{\text{space}}(f,g)$. We treat the second term of $\psi_{\text{space}}$, i.e., $\xi_{1}B_{1,1}(\eta)\partial_{\eta_{1}}\phi$ only since the other terms can be treated in a similar way. Let
$I_{\text{space,2}}(f,g):=\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}B_{1,1}(\eta)\partial_{\eta_{1}}\phi(\xi,\eta)e^{-i\tau\phi(\xi,\eta)}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau.$
By an integration by parts in $\eta_{1}$, we have
$I_{\text{space,2}}(f,g)$ $=$ $i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}\partial_{\eta_{1}}\left\{e^{-i\tau\phi(\xi,\eta)}\right\}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$=$ $-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}\partial_{\eta_{1}}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$+i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}B_{1,1}(\eta)\partial_{\eta_{1}}\widehat{f}(\xi-\eta)\widehat{g}(\eta)d\eta d\tau$
$-i\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}e^{-i\tau\phi(\xi,\eta)}B_{1,1}(\eta)\widehat{f}(\xi-\eta)\partial_{\eta_{1}}\widehat{g}(\eta)d\eta d\tau.$